OpenStudy (anonymous):

A cylinder labeled X has a height of 12 inches and a radius of 10 inches. Another cylinder labeled Y has a height of 36 inches and a radius of 30 inches. The volume of cylinder Y is _______ times the volume of cylinder X.

sam (.sam.):

Comparing both ratios, \[\large V_x:V_y\] \[\large\pi r^2_1h_1:\pi r^2_2h_2\] \[\large \cancel\pi r^2_1h_1:\cancel\pi r^2_2h_2\] \[\large r^2_1h_1: r^2_2h_2\] \[\large (10)^2(12): (30)^2(36)\] \[\large 1200:32400\] Do the final step by dividing 1200 to both sides

OpenStudy (anonymous):

1:27?

sam (.sam.):

yes